OpenStudy (anonymous):

Find the intervals of increase and decrease for the following functions. a) h(x)=x^3(x-1)^4 b)\[y=x \sqrt{4-x}\]

myininaya (myininaya):

Have you found dh/dx

OpenStudy (anonymous):

I tried and I got this dh/dt= x^2(4x^2-x-3). I'm not sure if it's right...

myininaya (myininaya):

i don't know if that is it or not i will check in a moment you did use product rule then power rule and chain rule right?

OpenStudy (anonymous):

Yes

myininaya (myininaya):

\[h=x^3 (x-1)^4 \\ \frac{dh}{dx}=(x-1)^4 \cdot \frac{d(x^3)}{dx}+x^3 \cdot \frac{d(x-1)^4}{dx}\]

myininaya (myininaya):

so you differentiated x^3 (w.r.t. x) and got ? so you differentiated (x-1)^4 (w.r.t. x) and got ?

OpenStudy (anonymous):

3x^2 4(x-1)^3

myininaya (myininaya):

\[\frac{dh}{dx}=(x-1)^4 \cdot 3x^2+x^3 \cdot 4(x-1)^3\] right

myininaya (myininaya):

now we can factor out the x^2(x-1)^3

myininaya (myininaya):

\[\frac{dh}{dx}=x^2(x-1)^3[3(x-1)+4x]\]

myininaya (myininaya):

now you need to find when dh/dx=0

OpenStudy (anonymous):

So x= 3/7 and 0?

myininaya (myininaya):

x^2=0 when x=0 x-1=0 when x=? 3(x-1)+4x=0 when x=3/7 2 out 3 correct

OpenStudy (anonymous):

oops! The last one is x=1

myininaya (myininaya):

ok cool so those are our critical numbers

myininaya (myininaya):

 we have 4 intervals to test

myininaya (myininaya):

(-inf,0) (0,3/7) (3/7,1) (1,inf)

myininaya (myininaya):

select a number from each then dh/dx for each of the numbers you picked (pick only 4 one from each interval )

myininaya (myininaya):

if dh/dx>0 then h is increasing if dh/dx<0 then h is decreasing

OpenStudy (anonymous):

So the intervals are : increases on (-inf,3/7) and (1,inf) decreases on (3/7,1) ???

myininaya (myininaya):

So a number from (-inf,0) is -1 \[\frac{dh}{dx}=x^2(x-1)^3(7x-3) \\ \frac{dh}{dx}|_{x=-1}=(-1)^2(-1-1)^3(7 \cdot -1 -3)=1(-2)^3(-7-3) \\ =-8(-10)=80>0 \text{ so } h \text{ is increasing on } (-\inf,0)\] now you tested also (0,3/7)? choose an ugly number like 2/7

myininaya (myininaya):

\[\frac{dh}{dx}|_{x=\frac{2}{7}}=(\frac{2}{7})^2(\frac{2}{7}-1)^3(7 \cdot \frac{2}{7}-3) \\ =pos \cdot negative \cdot negative=pos \]

myininaya (myininaya):

so i would say it is increasing on (-inf,0) U (0,3/7)

OpenStudy (anonymous):

oh okay!

myininaya (myininaya):

yep it looks good except I would separate that one interval you have

myininaya (myininaya):

increases on (-inf,0) or (0,3/7) or (1,inf) decreases on (3/7,1)